S10=-20 and d=4 find a1

Answer by stanbon(75887)

S10=-20 and d=4 find a1
-----
S(10) = a(1) + (n-1)d
-20 = a(1) + 9*4
a(1) = -20-36
a(1) = -56
================
Cheers,
Stan H.
==================